A)

The polynomial g(X)=3x^3-2x^2-12x+8


(I) using the factor theorem, show that (X+2) is a factor of g(X)


(II) factorize g(X) completely


1
Expert's answer
2021-09-28T14:42:22-0400

(i)


"x+2=0=>x=-2"

"g(-2)=3(-2)^3-2(-2)^2-12(-2)+8"

"=-24-8+24+8=0"

Since "g(-2)=0," then by thefactor theorem "(x+2)" is a factor of "g(x)."


(ii)


"g(x)=3x^3-2x^2-12x+8"

"=3x^2(x+2)-8x(x+2)+4(x+2)"

"=(x+2)(3x^2-8x+4)"

"=(x+2)(3x(x-2)-2(x-2))"

"=(x+2)(x-2)(3x-2)"